# Rui Lia,  Anyao Wangb   Mingqing Zhaic

### v2

@@ 4 Extremal graphs
With Claim 4.1 established, we can classify the components of $G[U]$ into three distinct categories:
  • $\mathcal{N}_{1}$ includes all components that are not trees.
  • $\mathcal{N}_{2}$ contains all components that are nontrivial trees.
  • $\mathcal{N}_{3}$ consists solely of isolated vertices.
Moreover, we denote
$g_{0}=t|V|+r-e(V),$
and for a graph $H$ with $V(H)\subseteq U$, we define
$\eta(H)=e(H)+(\lfloor t\rfloor-t)|V(H)|.$
Let $F\subseteq G[U]$ and $H$ is a component of $F$, we denote that
$\eta^{+}(F)=\sum_{\eta(H)>0}\eta(H).$
Obviously, $g_{0}$ is nonnegative and constant as $r\geq-\binom{\lfloor t+1\rfloor}{2}$ and $V$ is a $\lfloor t\rfloor$-clique. By the definitions of $g_{0}$ and $\eta(H)$, we immediately obtain the following observation.
Fact 4.2.
If $K_{|V|}\nabla H$ satisfies the property $P_{t,r}$, then $\eta(H)\leq g_{0}$. Moreover, if $F$ is a forest with $\eta^{+}(F)\leq g_{0}$, then $K_{|V|}\nabla F$ satisfies the property $P_{t,r}$.
Given that each component $H$ in $\mathcal{N}_{1}$ has at least $|V(H)|$ edges, we have $e(H)\geq|V(H)|$. Thus, for any such component $H$:
$\eta(H)=e(H)+\lfloor t\rfloor|V(H)|-t|V(H)|\geq|V(H)|+\lfloor t\rfloor|V(H)|-t|V(H)|\geq t_{0}|V(H)|.$
Since $t_{0}>0$, it follows that $\eta(H)>0$. Summing over all components $H\in\mathcal{N}_{1}$, we have
$t_{0}\sum_{H\in\mathcal{N}_{1}}|V(H)|\leq\sum_{H\in\mathcal{N}_{1}}\eta(H)\leq g_{0}.$
Therefore,
$\sum_{H\in\mathcal{N}_{1}}|V(H)|\leq\frac{g_{0}}{t_{0}}.$
This inequality indicates that the total number of vertices across all components in $\mathcal{N}_{1}$ is bounded by $\frac{g_{0}}{t_{0}}$.
For a component $H$ in $\mathcal{N}_{2}$, we have $e(H)=|V(H)|-1$. Thus, the function $\eta(H)$ for such a component becomes
$\eta(H)=(|V(H)|-1)+\lfloor t\rfloor|V(H)|-t|V(H)|=t_{0}|V(H)|-1.$
For $\eta(H)>0$,
$t_{0}|V(H)|-1>0\implies|V(H)|>\frac{1}{t_{0}}.$
This implies that components with $\eta(H)>0$ must have more than $\frac{1}{t_{0}}$ vertices. And similar to above argument, the total number of vertices of these components is bounded by $\frac{g_{0}+1}{t_{0}}$. Conversely, if $\eta(H)\leq 0$, then
$t_{0}|V(H)|-1\leq 0\implies|V(H)|\leq\frac{1}{t_{0}}.$
Such components have a limited number of vertices, which is at most $\lfloor\frac{1}{t_{0}}\rfloor$. Therefore, the remaining components, having $\eta(H)\leq 0$, contain lots of vertices that is $\Theta(n)$.
@@
Each component $H$ in $\mathcal{N}_{2}$ is a star.
@@
Let $H\in\mathcal{N}_{2}$ be a tree. Then $e(H)=|V(H)|-1$. According to Lemma 4.3, the star is walk-preferable to other trees with the same size. And by Lemma 4.1, a graph which has a walk-preferable subgraph in a same vertex set has a greater spectral radius. Since $G$ has a maximal spectral radius, $H$ must be a star. ∎
Having characterized the components in $\mathcal{N}_{2}$ as stars, we next examine the family $\mathcal{N}_{1}$.
@@
The family $\mathcal{N}_{1}$ contains at most one component.
@@
We aim to show that the family $\mathcal{N}_{1}$ contains at most one component. To do this, we will proceed by contradiction. Suppose there are two distinct components $H_{1}$ and $H_{2}$ in $\mathcal{N}_{1}$. Since each component in $\mathcal{N}_{1}$ contains at least one cycle, both $H_{1}$ and $H_{2}$ must have at least as many edges as vertices. Therefore, we get
$e(H_{1})+e(H_{2})\geq|V(H_{1})|+|V(H_{2})|\geq 6.$
Define $a=e(H_{1})+e(H_{2})-|V(H_{1})|-|V(H_{2})|+1$. Now, we consider the graph $H$ formed by the union of $H_{1}$, $H_{2}$, and $a$ copies of the star $K_{1,\lfloor\frac{1}{t_{0}}\rfloor-1}$ (note that there are sufficient many such stars.):
$H=H_{1}\cup H_{2}\cup aK_{1,\lfloor\frac{1}{t_{0}}\rfloor-1}.$
If $\lfloor\frac{1}{t_{0}}\rfloor=1$, then $H$ simplifies to
$H=H_{1}\cup H_{2}\cup aK_{1}.$
In either case, it is clear that
$e(H)=|V(H)|-1.$
Transform $H$ into a star $H^{\prime}=K_{1,|V(H)|-1}$. Starting from the original graph $G$, construct a new graph $G^{\prime}$ by removing the edges of $H$ and adding the edges of $H^{\prime}$. Note that $e(H)=e(H^{\prime})$ and $V(H)=V(H^{\prime})$. The parameter $\eta$ for $H^{\prime}$ is calculated as follows:
$\eta(H^{\prime})=e(H^{\prime})-(1-t_{0})|V(H^{\prime})|.$
Substituting the values:
$\eta(H^{\prime})$ $=e(H_{1})+e(H_{2})+a(\lfloor\frac{1}{t_{0}}\rfloor-1)-(1-t_{0})(|V(H_{1})|+|V(H_{2})|+a\lfloor\frac{1}{t_{0}}\rfloor)$
$=e(H_{1})-(1-t_{0})|V(H_{1})|+e(H_{2})-(1-t_{0})|V(H_{2})|-a+at_{0}\lfloor\frac{1}{t_{0}}\rfloor$
$=\eta(H_{1})+\eta(H_{2})-a(1-t_{0}\lfloor\frac{1}{t_{0}}\rfloor).$
Since $1-t_{0}\lfloor\frac{1}{t_{0}}\rfloor\geq 0$ and $a\geq 0$, it follows that
$\eta(H^{\prime})\leq\eta(H_{1})+\eta(H_{2})\leq g_{0}.$
By Fact 4.2, $G^{\prime}$ satisfies the property $P_{t,r}$. By Lemma 4.3, we have $F_{2}(H)<F_{2}(H^{\prime})$. According to Fact 4.1, we get $\rho(G)<\rho(G^{\prime})$. This contradicts the assumption that $G$ was optimal (or maximal). Therefore, $\mathcal{N}_{1}$ can contain at most one component. ∎
Having constrained the number of non-tree components in $\mathcal{N}_{1}$, we now turn our attention to the tree components in $\mathcal{N}_{2}$, particularly those that are stars. We aim to show that there can be at most one star of a specific size in $\mathcal{N}_{2}$.
@@
$\mathcal{N}_{2}$ has at most one star $K_{1,\lfloor\frac{1}{t_{0}}\rfloor+a}$ with $a\geq 1$.
@@
Suppose, for contradiction, that there are two stars $K_{1,\lfloor\frac{1}{t_{0}}\rfloor+n_{1}}$ and $K_{1,\lfloor\frac{1}{t_{0}}\rfloor+n_{2}}$ in $\mathcal{N}_{2}$ with $n_{1}\geq n_{2}\geq 1$. Construct a new graph $G^{\prime}$ from $G$ by a edge-shifting operation, i.e., replacing $K_{1,\lfloor\frac{1}{t_{0}}\rfloor+n_{1}}\cup K_{1,\lfloor\frac{1}{t_{0}}\rfloor+n_{2}}$ with $K_{1,\lfloor\frac{1}{t_{0}}\rfloor+n_{1}+1}\cup K_{1,\lfloor\frac{1}{t_{0}}\rfloor+n_{2}-1}$. Note that
$\eta(K_{1,\lfloor\frac{1}{t_{0}}\rfloor+n_{1}+1})+\eta(K_{1,\lfloor\frac{1}{t_{0}}\rfloor+n_{2}-1})=\eta(K_{1,\lfloor\frac{1}{t_{0}}\rfloor+n_{1}})+\eta(K_{1,\lfloor\frac{1}{t_{0}}\rfloor+n_{2}})\leq g_{0},$
and
$F_{2}(K_{1,\lfloor\frac{1}{t_{0}}\rfloor+n_{1}}\cup K_{1,\lfloor\frac{1}{t_{0}}\rfloor+n_{2}})<F_{2}(K_{1,\lfloor\frac{1}{t_{0}}\rfloor+n_{1}+1}\cup K_{1,\lfloor\frac{1}{t_{0}}\rfloor+n_{2}-1}).$
Since $(\lfloor\frac{1}{t_{0}}\rfloor+1)t_{0}-1>\frac{1}{t_{0}}t_{0}-1=0$ and $n_{2}\geq 1$, we have
$g_{0}\geq\eta(K_{1,\lfloor\frac{1}{t_{0}}\rfloor+n_{2}-1})=(\lfloor\frac{1}{t_{0}}\rfloor+n_{2})t_{0}-1=(\lfloor\frac{1}{t_{0}}\rfloor+1)t_{0}-1+(n_{2}-1)t_{0}>0.$
It follows that
$\eta(K_{1,\lfloor\frac{1}{t_{0}}\rfloor+n_{1}+1})\leq g_{0}-\eta(K_{1,\lfloor\frac{1}{t_{0}}\rfloor+n_{2}-1})<g_{0}.$
Consequently, $G^{\prime}$ satisfies the property $P_{t,r}$ and $\rho(G^{\prime})>\rho(G)$, a contradiction. This completes the proof. ∎
@@
With the constraints on $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ established, we can now link these findings to the overall structure of $G[U]$. Specifically, we investigate the implications of having a single non-tree component in $\mathcal{N}_{1}$.
@@
  1. (1)
    $|\mathcal{N}_{1}|=1$.
  2. (2)
    $43t_{0}\leq g_{0}<8t_{0}-$ 4.
  3. (3)
    $G[U]\cong K_{3}\cup(n-\lfloor t\rfloor-3)K_{1}$.
@@
$(1)\Rightarrow(2)$: Assume $\mathcal{N}_{1}=\{H_{1}\}$, meaning there is exactly one non-tree component in $G[U]$, which implies $|V(H_{1})|\geq 3$. We will show that $H_{1}\cong K_{3}$.
If $e(H_{1})\geq 4$, we proceed similarly to the proof of Claim 4.3. Consider replacing $H:=H_{1}\cup aK_{1,\lfloor\frac{1}{t_{0}}\rfloor-1}$ with a star $H^{\prime}:=K_{1,|V(H)|-1}$, where $a=e(H_{1})+1-|V(H_{1})|$. We can show that this replacement satisfies the property $P_{t,r}$. By Lemmas 4.1 and 4.3, this replacement would yield a new graph with a greater spectral radius than $G$, leading to a contradiction. Therefore, $e(H_{1})=3$.
Given $e(H_{1})=3$ and $|V(H_{1})|\geq 3$, $H_{1}$ must be either $K_{3}$ or $K_{1,3}$ (a claw). Since $W_{3}(K_{3})=24>W_{3}(K_{1,3})=18$, $K_{3}$ is walk-preferable to the star $K_{1,3}$. Thus, $H_{1}\cong K_{3}$.
It follows that $\eta(H_{1})\geq 3t_{0}$, confirming that $H_{1}$ is indeed a triangle.
Next, consider $\eta(4K_{2})=8t_{0}-4$. If $g_{0}\geq 8t_{0}-4$, then $e(U)\geq 4$, as $4K_{2}$ is walk-preferable to $K_{3}\cup 5K_{1}$. This would imply the existence of a star $K_{1,a}$ in $\mathcal{N}_{2}$ with $a\geq 1$. Replacing $K_{3}\cup K_{1,a}$ with $K_{1,3+a}$ would again result in a graph with a greater spectral radius, contradicting our assumption. Therefore, $g_{0}<8t_{0}-4$.
$(2)\Rightarrow(3)$: Given $43t_{0}\leq g_{0}<8t_{0}-$ 4, the upper bound on $g_{0}$ restricts the number of edges in $U$. Specifically, since $\eta(H)\geq 8t_{0}-4$ for any graph $H$ with $e(H)=4$, it follows that $e(U)\leq 3$. Moreover, because $\eta(K_{3})=3t_{0}\leq g_{0}$, we conclude that $e(U)=3$. By Lemma 4.1 and previous discussions, the only graph that satisfies these conditions is $G[U]\cong K_{3}\cup(n-\lfloor t\rfloor-3)K_{1}$.
$(3)\Rightarrow(1)$: If $G[U]\cong K_{3}\cup(n-\lfloor t\rfloor-3)K_{1}$, then $G[U]$ contains exactly one non-tree component $K_{3}$. Hence, $|\mathcal{N}_{1}|=1$.
This completes the proof of the claim. ∎
Based on the preceding analysis, we can now establish the following theorem, which characterizes the structure of the extremal graph $G$ that satisfies the property $P_{t,r}$ under the condition $43t_{0}\leq g_{0}<8t_{0}-$ 4.
@@
Let $G$ be the spectral extremal graph that satisfies the property $P_{t,r}$. If $43t_{0}\leq g_{0}<8t_{0}-$ 4, then $G$ is isomorphic to the join of a clique $K_{\lfloor t\rfloor}$ and a graph $H$, where $H\cong K_{3}\cup(n-\lfloor t\rfloor-3)K_{1}$. In other words,

### v3

@@ 4 Characterization of extremal graphs
By Lemma 4.5 and Claim 4.1, we have $\eta^{+}(G[U])\leq g_{0}$. Next, we classify the components of $G[U]$ into three distinct categories $\mathcal{N}_{1},$ $\mathcal{N}_{2}$ and $\mathcal{N}_{3}$, where
(i) $\mathcal{N}_{1}$ includes all components that are not trees;
(ii) $\mathcal{N}_{2}$ contains all tree components of order at least $\lfloor 1/t_{0}\rfloor+1$;
(iii) $\mathcal{N}_{3}$ consists of all tree components of order at most $\lfloor 1/t_{0}\rfloor$.
@@
Every component in $\mathcal{N}_{1}$ is positive. Moreover, $\sum_{F\in\mathcal{N}_{1}}|V(F)|\leq g_{0}/t_{0}.$
@@
For each $F\in\mathcal{N}_{1}$, we have $e(F)\geq|V(F)|$. Recall that $t_{0}=\lfloor t\rfloor+1-t>0$. Thus,
$\eta(F)=e(F)+(\lfloor t\rfloor-t)|V(F)|\geq t_{0}|V(F)|>0.$ (16)
Hence, $F$ is a positive component. Furthermore, summing over all components $F\in\mathcal{N}_{1}$, we obtain that $\sum_{F\in\mathcal{N}_{1}}\eta(F)\geq t_{0}\sum_{F\in\mathcal{N}_{1}}|V(F)|$. However, by Lemma 4.5, we know that $\sum_{F\in\mathcal{N}_{1}}\eta(F)\leq\eta^{+}(G[U])\leq g_{0}$ since $G\cong K_{\lfloor t\rfloor}\nabla G[U]$ and $G$ satisfies the property $P_{t,r}$. Therefore, we have $\sum_{F\in\mathcal{N}_{1}}|V(F)|\leq g_{0}/t_{0},$ as claimed. ∎
@@
Every component in $\mathcal{N}_{2}$ is positive and $\sum_{F\in\mathcal{N}_{2}}|V(F)|$ is constant.
@@
For each $F\in\mathcal{N}_{2}$, we have $e(F)=|V(F)|-1$ and $|V(F)|\geq\lfloor 1/t_{0}\rfloor+1$. Thus,
$\eta(F)=e(F)+(\lfloor t\rfloor-t)|V(F)|=t_{0}|V(F)|-1>0.$ (17)
Hence, all components in $\mathcal{N}_{2}$ are positive and $\sum_{F\in\mathcal{N}_{2}}\eta(F)=t_{0}\sum_{F\in\mathcal{N}_{2}}|V(F)|-|\mathcal{N}_{2}|$, where $\sum_{F\in\mathcal{N}_{2}}|V(F)|\geq|\mathcal{N}_{2}|(\lfloor 1/t_{0}\rfloor+1)$. Thus,
$\sum_{F\in\mathcal{N}_{2}}\eta(F)\geq\Big{(}t_{0}-\frac{1}{\lfloor 1/t_{0}\rfloor+1}\Big{)}\sum_{F\in\mathcal{N}_{2}}|V(F)|.$
On the other hand, by Lemma 4.5, we have $\sum_{F\in\mathcal{N}_{2}}\eta(F)\leq\eta^{+}(G[U])\leq g_{0}$. Therefore, $\sum_{F\in\mathcal{N}_{2}}|V(F)|$ is upper bounded by $g_{0}/\big{(}t_{0}-\frac{1}{\lfloor 1/t_{0}\rfloor+1}\big{)}.$
@@
Every component in $\mathcal{N}_{2}\cup\mathcal{N}_{3}$ is a star.
@@
Suppose to the contrary that $\mathcal{N}_{2}\cup\mathcal{N}_{3}$ contains some components that are not stars. Then, let $G^{\prime}$ be the graph obtained from $G$ by replacing each component in $\mathcal{N}_{2}\cup\mathcal{N}_{3}$ with a star of the same order. By Lemma 4.3, the star is strictly walk-preferable to other trees with the same order. By Claim 4.3 and the definition of $\mathcal{N}_{3}$, every component in $\mathcal{N}_{2}\cup\mathcal{N}_{3}$ has finite order. Hence, by Corollary 4.1, we have $\rho(G^{\prime})>\rho(G)$.
However, in view of (17), we know that $\eta(F)=t_{0}|V(F)|-1$ for each tree $F\in\mathcal{N}_{2}\cup\mathcal{N}_{3}$. This implies that $\eta^{+}(G^{\prime}[U])=\eta^{+}(G[U])\leq g_{0}$. By Lemma 4.5, $G^{\prime}$ satisfies the property $P_{t,r}$, which leads to a contradiction. ∎
From (15), it follows that $\eta(F)\leq 0$ for every tree $F$ with $|V(F)|\leq\lfloor 1/t_{0}\rfloor$. Hence, no component in $\mathcal{N}_{3}$ is positive. Furthermore, the following statement holds for $\mathcal{N}_{3}$.
Claim 4.5.
There is at most one star $F$ in $\mathcal{N}_{3}$ with $|V(F)|\leq\lfloor 1/t_{0}\rfloor-1$.
Proof.
@@
Suppose, for contradiction, that there are two stars $K_{1,a}$ and $K_{1,b}$ in $\mathcal{N}_{3}$ such that $20\leq b\leq a\leq\lfloor 1/t_{0}\rfloor-$ 2. If $b\geq 1$, we construct a graph $G^{\prime}$ from $G$ by performing an edge-shifting operation, i.e., replacing $K_{1,a}\cup K_{1,b}$ with $K_{1,a+1}\cup K_{1,b-1}$. In light of (15), neither $K_{1,a+1}$ nor $K_{1,b-1}$ are positive. Thus, $\eta^{+}(G^{\prime}[U])=\eta^{+}(G[U])\leq g_{0}$. By Lemma 4.5, $G^{\prime}$ satisfies the property $P_{t,r}$. Notice that $F_{2}(K_{1,a+1}\cup K_{1,b-1})>F_{2}(K_{1,a}\cup K_{1,b})$. By Corollary 4.1, we have $\rho(G^{\prime})>\rho(G)$, which contradicts the maximality of $\rho(G)$.
If $b=0$, i.e., $K_{1,b}$ is an isolated vertex, we construct a new graph $G^{\prime\prime}$ from $G$ by adding an edge between the center of $K_{1,a}$ and the isolated vertex. It is clear that $\rho(G^{\prime\prime})>\rho(G)$. Moreover, $\eta(K_{1,a+1})\leq 0$, and hence $\eta^{+}(G^{\prime\prime}[U])=\eta^{+}(G[U])\leq g_{0}$. Therefore, $G^{\prime\prime}$ satisfies the property $P_{t,r}$, which also leads to a contradiction. ∎
If $\mathcal{N}_{2}$ contains two stars of order at least $\lfloor 1/t_{0}\rfloor+2$, we can also perform an edge-shifting operation, i.e., replacing $K_{1,\lfloor 1/t_{0}\rfloor+a}\cup K_{1,\lfloor 1/t_{0}\rfloor+b}$ with $K_{1,\lfloor 1/t_{0}\rfloor+a+1}\cup K_{1,\lfloor 1/t_{0}\rfloor+b-1}$. Using a similar argument as in Claim 4.5, we obtain the following statement.
Claim 4.6.
There is at most one star $K_{1,\lfloor\frac{1}{t_{0}}\rfloor+a}$ in $\mathcal{N}_{2}$ with $a\geq 1$.
By Claims 4.2 and 4.3, $\sum_{F\in\mathcal{N}_{3}}|V(F)|=\Theta(n)$. Furthermore, by Claim 4.5, $\mathcal{N}_{3}$ contains $\Theta(n)$ copies of $K_{1,\lfloor 1/{t_{0}}\rfloor-1}$. Based on this, we have the following assertion.
Claim 4.7.
There is at most one component $F$ in $\mathcal{N}_{1}$. If $|\mathcal{N}_{1}|=1$, then $F$ is a triangle.
Proof.
Suppose to the contrary that there are two components $F_{1},F_{2}\in\mathcal{N}_{1}$. Then $e(F_{i})\geq|V(F_{i})|\geq 3$ for $i\in\{1,2\}$. Let $\sum_{i=1}^{2}e(F_{i})=\sum_{i=1}^{2}|V(F_{i})|-1+a$. By Claim 4.2, we obtain $\sum_{i=1}^{2}|V(F_{i})|\leq g_{0}/t_{0}$, and thus $a$ is also bounded by a constant. Recall that $\mathcal{N}_{3}$ contains $\Theta(n)$ stars of order $\lfloor 1/t_{0}\rfloor$. Now, consider the subgraph $F$ formed by the union of $F_{1}$, $F_{2}$, and $a$ stars $K_{1,\lfloor 1/t_{0}\rfloor-1}$ from $\mathcal{N}_{3}$. It is clear that $e(F)=|V(F)|-1$, and $|V(F)|$ is finite. In view of (15), we have $\eta(K_{1,\lfloor 1/t_{0}\rfloor-1})\leq 0$. Recall that $t_{0}=1+\lfloor t\rfloor-t$. Then
$\eta^{+}(F)=\sum_{i=1}^{2}\eta(F_{i})=\sum_{i=1}^{2}e(F_{i})+(\lfloor t\rfloor-t)\sum_{i=1}^{2}|V(F_{i})|=t_{0}\sum_{i=1}^{2}|V(F_{i})|-1+a.$ (18)
Now, we transform $F$ into a star $F^{\prime}=K_{1,|V(F)|-1}$ and construct a graph $G^{\prime}$ from $G$ by replacing the edges of $F$ with the edges of $F^{\prime}$. From (17), we observe that $\eta^{+}(F^{\prime})=\eta(F^{\prime})=t_{0}|V(F^{\prime})|-1$, where $|V(F^{\prime})|=\sum_{i=1}^{2}|V(F_{i})|+a\lfloor{1}/{t_{0}}\rfloor$. By comparing with (18), it is straightforward to verify that $\eta^{+}(F^{\prime})\leq\eta^{+}(F)$. Thus, $\eta^{+}(G^{\prime}[U])\leq\eta^{+}(G[U])\leq g_{0}$, and hence $G^{\prime}$ satisfies the property $P_{t,r}$. Notice that $e(F^{\prime})=e(F)>3$. By Lemma 4.3, we see that $F_{2}(F^{\prime})>F_{2}(F)$. Since $|V(F)|$ is finite, by Corollary 4.1, we obtain $\rho(G^{\prime})>\rho(G)$, which leads to a contradiction. Therefore, $|\mathcal{N}_{1}|\leq 1$.
Next, assume that $|\mathcal{N}_{1}|=\{F_{1}\}$. We will demonstrate that $F_{1}$ is a triangle. The proof follows similarly to the previous one. If $e(F_{1})\geq 4$, we denote $e(F_{1})=|V(F_{1})|-1+a$ and $F=F_{1}\cup aK_{1,\lfloor 1/t_{0}\rfloor-1}$. Let $G^{\prime}$ be a graph obtained from $G$ by removing the edges of $F$ and adding the edges of $F^{\prime}$, where $F^{\prime}=K_{1,|V(F)|-1}$. We also have $\eta^{+}(F^{\prime})\leq\eta^{+}(F),$ and hence $\eta^{+}(G^{\prime}[U])\leq\eta^{+}(G[U])\leq g_{0}$, which implies that $G^{\prime}$ satisfies the property $P_{t,r}$. By a similar argument as above, we can conclude that $\rho(G^{\prime})>\rho(G)$, a contradiction. Therefore, $e(F_{1})=3$, and so $F_{1}$ is a triangle. This completes the proof. ∎
@@
With the constraints on $\mathcal{N}_{1},\mathcal{N}_{2}$ and $\mathcal{N}_{3}$ established as outlined above, we can now relate these findings to the overall structure of $G[U]$. Specifically, we will investigate the implications of having a single component in $\mathcal{N}_{1}$.
@@
(i) $|\mathcal{N}_{1}|=1$;
(ii) $43t_{0}\leq g_{0}<8t_{0}-$ 4;
(iii) $G[U]\cong K_{3}\cup(n-\lfloor t\rfloor-3)K_{1}$.
@@
We prove the equivalence of the three statements by showing the implications $(i)\Rightarrow(ii)$, $(ii)\Rightarrow(iii)$, and $(iii)\Rightarrow(i)$.
$(i)\Rightarrow(ii)$: Assume that $|\mathcal{N}_{1}|=1$. By Claim 4.7, the unique component in $\mathcal{N}_{1}$ is a triangle $K_{3}$. Recall that $t_{0}=\lfloor t\rfloor+1-t>0$. Then $\eta(K_{3})=e(K_{3})+(\lfloor t\rfloor-t)|V(K_{3})|=3t_{0}$. Since $\eta(K_{3})\leq\eta^{+}(G[U])\leq g_{0}$, we have $g_{0}\geq 3t_{0}$.
Next, we show that $g_{0}<8t_{0}-4$. Suppose, for contradiction, that $g_{0}\geq 8t_{0}-4$. Since $\mathcal{N}_{1}$ contains a triangle, we have $e(G[U])\geq 4$. Otherwise, $\mathcal{N}_{2}\cup\mathcal{N}_{3}$ would consist only of isolated vertices. Recall that $\mathcal{N}_{3}$ contains $\Theta(n)$ components. We can construct a new graph $G^{\prime}$ from $G$ by replacing $K_{3}\cup 5K_{1}$ with $4K_{2}$. We observe that $\eta^{+}(G^{\prime}[U])\leq\eta(4K_{2})=8t_{0}-4\leq g_{0}$. Thus, by Lemma 4.5, $G^{\prime}$ satisfies the property $P_{t,r}$. However, since $4K_{2}$ is walk-preferable to $K_{3}\cup 5K_{1}$, Corollary 4.1 implies that $\rho(G^{\prime})>\rho(G)$, a contradiction. This contradiction establishes the existence of a star $K_{1,a}$ in $\mathcal{N}_{2}\cup\mathcal{N}_{3}$ with $a\geq 1$.
Now, we construct another graph $G^{\prime\prime}$ from $G$ by replacing $K_{3}\cup K_{1,a}$ with $K_{1,3+a}$. By Lemma 4.3 and Corollary 4.1, we have $\rho(G^{\prime\prime})>\rho(G)$. Since $K_{1,3+a}$ and $K_{3}\cup K_{1,a}$ have the same number of vertices and the same number of edges, we have $\eta(K_{1,3+a})=\eta(K_{3}\cup K_{1,a})$, and hence, $\eta^{+}(G^{\prime\prime}[U])=\eta^{+}(G[U])\leq g_{0}.$ Thus, $G^{\prime\prime}$ also satisfies the property $P_{t,r}$, contradicting the maximality of $\rho(G)$. Therefore, $g_{0}<8t_{0}-4$.
$(ii)\Rightarrow(iii)$: Assume that $43t_{0}\leq g_{0}<8t_{0}-$ 4. Then $t_{0}>\frac{4}{5}$. If $e(G[U])\leq 2$, we modify the graph $G$ by removing the edges within $U$ and then embedding a triangle in $U$. By a similar argument as above, we get a contradiction. Thus, $e(G[U])\geq 3$.
Let $F$ be the subgraph induced by all the edges in $G[U]$. Clearly, $|V(F)|\leq 2e(F)$, and thus, $\eta(F)=e(F)-(1-t_{0})|V(F)|\geq e(F)(2t_{0}-1).$ If $e(F)\geq 4$, then $\eta(F)\geq 8t_{0}-4$ and so $\eta^{+}(G[U])=\eta(F)>g_{0}$. This is a contradiction. Thus, we have $e(F)=3$. Furthermore, by Lemma 4.3, $F$ must be either $K_{1,3}$ or $K_{3}$. Since $W_{3}(K_{3})=24>W_{3}(K_{1,3})=18$, $K_{3}$ is walk-preferable to $K_{1,3}$. Therefore, $F\cong K_{3}$, and so $G[U]\cong K_{3}\cup(n-\lfloor t\rfloor-3)K_{1}$.
$(iii)\Rightarrow(i)$: If $G[U]\cong K_{3}\cup(n-\lfloor t\rfloor-3)K_{1}$, then $G[U]$ contains exactly one non-tree component $K_{3}$. Hence, $|\mathcal{N}_{1}|=1$.
This completes the proof. ∎
Building on the preceding analysis and Claim 4.8, we can now establish the following theorem, which uniquely determines the extremal graph $G$ when $43t_{0}\leq g_{0}<8t_{0}-$ 4.
@@
Let $G$ be a graph that attains the maximum spectral radius over all $n$-vertex graphs satisfying the property $P_{t,r}$. If $43t_{0}\leq g_{0}<8t_{0}-$ 4, then we have